The internal energy of a system is 513 J and the quantity of work done by the system is 50 J. How much heat is absorbed?


1
Expert's answer
2022-05-23T16:49:14-0400
"E=U+W""E=513J+50J=563J"

The system absorbs 563 J of heat.